OpenStudy (anonymous):

Which formula gives the coordinates of the midpoint of the segment connecting points (a, b) and (c, d)?

OpenStudy (anonymous):

@KamiBug

OpenStudy (solomonzelman):

\(\large\color{black}{ \left(\begin{matrix} \frac{\LARGE a+c }{\LARGE 2}~,~\frac{\LARGE b+d }{\LARGE 2} \\ \end{matrix}\right) }\) \(\large\color{black}{ \Uparrow~~~~~~~~~\Uparrow }\) \(\large\color{black}{ \left(\begin{matrix} ~x~~~~~,~~~~~y ~ \\ \end{matrix}\right) }\)

OpenStudy (solomonzelman):

you can ignore the arrows and the (x,y)

OpenStudy (anonymous):

these are the choices

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

you can't type them? (Sorry if this sounds rude)

OpenStudy (anonymous):

nope

OpenStudy (solomonzelman):

Hint: when you divide by 2, it is same as multiplying times 1/2.

OpenStudy (anonymous):

divide what?

OpenStudy (anonymous):

@SolomonZelman